Emily Hill is 24 years old and single, lives in an apartment, and has no dependents. Last year she earned $55,000 as a sales assistant for Business Solutions: $3,910 of her wages was withheld for federal income taxes. In addition, she had interest income of $142.

- Estimate her taxable income

- Estimate her tax liability

- Estimate her tax refund or tax owed
Answer #1

Answer:

Emily Hill is 24 years old and single and has no dependents.

Tax year = Last year = 2018

Standard deduction for single = $12,000

There is no personal exemption amount for 2018.

Taxable Income:

Earnings as sales assistant for Business Solutions = $55,000

Add,  interest income = $142

Gross Income = $55,000 + $142 = $55,142

Less, Standard deduction = $12,000

Taxable Income = 55142 - 12000 = $43,142

Taxable Income = $43,142

Tax Liability:

Single 2018 tax rates and brackets Taxable Income S0 - S9,525 $9,526- $38,700 Tax Rate 10% of taxable income $952.50 plus 12% of the amount over $9,525 $4.453.50 plus 22% of the amount over $38.700 $14,089.50 plus 24% of the amount over $82,500 $32,089.50 plus 32% of the amount over $157,500 $45,689.50 plus 35% of the amount over $200,000 $150,689.50 plus 37% of the amount over $500,000 $82,501- $157,500 $157,501- $200,000 $200,001 $500,000 $500,001 or more

Taxable Income = $43,142

Tax liability = $4453.50 + (43142 -38700) * 22% = $5431

Tax liability = $5,431

Tax owed

Tax owed = Tax liability - withholding = 5431 - 3910 = $1,521

Tax owed = $1,521